Bhim is 15 years older than his wife, 35 years older than his son, 37 years older than his daughter. bhim and his wife had been married for 5 years before the son was born. If their combined age added up to 167, how old was bhim's wife when they got married.

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)


let b = bhim's age today.
let w = bhim's wife's age today.
let s = bhim's son's age today.
let d = bhim's daughter's age today.

the sum of their age today is 167.

b + w + s + d = 167.

b = w + 15
b = s + 35
b = d + 37

solve for w,s,d to get:

w = b - 15
s = b - 35
d = b - 37

the equation of b + w + s + d = 167 becomes:

b + b - 15 + b - 35 + b = 37 = 167

combine like terms to get:

4b - 87 = 167

add 87 to both sides of this equation to get:

4b = 254

solve for b to get b = 254 / 4 = 63.5

bhim is 63.5 today.
bhim's wife is 63.5 - 15 = 48.5 today.
bhim's son is 63.5 - 35 = 28.5 today.
bhim's daughter is 63.5 - 37 = 26.5 today.

the sum of their ages is 63.5 + 48.5 + 28.5 + 26.5 = 167.

bhim was married 5 years before his son was born.

his son was born 28.5 years ago.

add 5 to that to get bhim was married 33.5 years ago.

33.5 years ago, bhim was 63.5 - 33.5 = 30.

his wife was 48.5 - 33.5 = 15.

that's your solution.

bhim's wife was 15 when she married him.


Answer by ikleyn(53763)

.
After reading the condition, it is not clear to me: whose combined age is 167 ?

Combined age of the four participants ?

Or Bhim and his wife only ?

If after reading the condition the reader has such questions, it means that the problem allows different interpretations -
and, therefore, is not a Math problem at all.